# Latent Ewald summation for machine learning of long-range interactions

### v1

@@ Abstract
Machine learning interatomic potentials (MLIPs) often neglect long-range interactions, such as electrostatic and dispersion forces. In this work, we introduce a straightforward and efficient method to account for long-range interactions by learning a latent variable from local atomic descriptors and applying an Ewald summation to this variable. We demonstrate that in systems including charged, polar, or apolar molecular dimers, bulk water, and water-vapor interface, standard short-ranged MLIPs can lead to unphysical predictions even when employing message passing. The long-range models effectively eliminate these artifacts, with only about twice the computational cost of short-range MLIPs.
I Introduction
@@
Machine learning interatomic potentials (MLIPs) can learn from reference quantum mechanical calculations and then predict the energy and forces of atomic configurations quickly, thus allowing for a more accurate and comprehensive exploration of material and molecular properties at scale Keith et al. (2021); Unke et al. (2021a). Most state-of-the-art MLIP methods use a short-range approximation: the effective potential energy surface experienced by one atom is determined by its atomic neighborhood. This approximation implies that the total energy is the sum of atomic contributions, which also makes the MLIPs scale linearly with system size.
@@
The short-range MLIPs, however, neglect all kinds of long-range interactions such as Coulomb and dispersion. Although short-range potentials may be sufficient for describing most properties of homogeneous bulk systems Yue et al. (2021), they may fail for liquid-vapor interfaces Niblett et al. (2021), dielectric response Rodgers and Weeks (2008); Cox (2020), dilute ionic solutions with Debye-Hückel screening, and interactions between gas phase molecules Huguenin-Dumittan et al. (2023).
@@
There has been a continuous effort to incorporate long-range interactions into MLIPs. One can include empirical electrostatics and dispersion baseline corrections Niblett et al. (2021); Lee et al. (2022); Unke et al. (2021b), but for many systems, such baseline is not readily available. Another option is to predict effective partial charges to each atom, which are then used to calculate long-range electrostatics Unke and Meuwly (2019); Ko et al. (2021); Gao and Remsing (2022); Sifain et al. (2018); Gong et al. (2024); Shaidu et al. (2024). For example, the fourth-generation high-dimensional neural network potential (4G-HDNNPs) Ko et al. (2021) predicts the electronegativities of each nucleus, and then use a charge equilibration scheme Rappe and Goddard III (1991) to assign the charges. 4G-HDNNPs are trained directly to reproduce atomic partial charges from reference quantum mechanical calculations, although partial charges are not physical observables and their values depend on the specific partitioning scheme used Sifain et al. (2018). In a similar vein, the deep potential long-range (DPLR) Zhang et al. (2022) learns maximally localized Wannier function centers (MLWFCs) for insulating systems, and the self-consistent field neural network (SCFNN) Gao and Remsing (2022) predicts the electronic response via the position of the MLWFCs. Message passing neural networks (MPNNs) Schütt et al. (2017); Batzner et al. (2022); Deng et al. (2023); Haghighatlari et al. (2022) employ a number of graph convolution layers to communicate information between atoms, thereby capturing long-range interaction up to the local cutoff radius times the number of layers. However, if parts of the system are disconnected on the graph, e.g. two molecules with a distance beyond the cutoff, the message passing scheme does not help. A very interesting approach is the long-distance equivariant (LODE) method Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023), which uses local descriptors to encode the Coulomb and other asymptotic decaying potentials ($1/r^{p}$) around the atoms.
@@ II Theory
where $E_{\theta}$ is a multilayer perceptron with parameters $\theta$ that maps the invariant features ($B$) of an atom to its short-range atomic energy. $B$ can be any invariant features used in different MLIP methods, including those based on local atomic environment descriptors such as ACE Drautz (2019), atom centered symmetry functions Behler and Parrinello (2007), smooth overlap of atomic positions (SOAP) Bartók et al. (2013), or any latent invariant features in MPNNs.
@@
where $\mathbf{k}=(2\pi n_{x}/L_{x},2\pi n_{y}/L_{y},2\pi n_{z}/L_{z})$ is a reciprocal vector of the orthorhombic cell, and $\mathbf{r}_{i}$ is the Cartesian coordinates of atom $i$. The long-range energy is then obtained using the Ewald summation:
@@
where $\sigma$ is a smearing factor which we typically set to 1 Å, $k=|\mathbf{k}|$ is the magnitude, and $k_{c}$ is the maximum cutoff. $q$ can be multi-dimensional, in which case the total long-range energy is aggregated over contributions from different dimensions of $q$ after the Ewald summation.
@@
The LES method can be interpreted in two ways. First, the hidden variable $q$ is analogous to environment-dependent partial charges on each atom. This implies that the method is at least as expressive as those explicitly based on learning partial atomic charges, as it would yield the same results if $q$ replicated the partial charges. Additionally, $q$ can be multi-dimensional, potentially enhancing expressiveness further. Unlike partial charges, $q$ is not constrained by requirements such as charge neutrality or correct charge magnitudes. Since the Ewald summation of $q$ doesn’t need to correspond to a physical electrostatic potential, Eqn. (4) omits the self-interaction term present in the regular Ewald summation for long-range charge interactions. The second interpretation of LES is as a mechanism that allows atoms far apart in the simulation box to communicate their local information. In this sense, LES is related to the recent Ewald-based long-range message-passing method, which facilitates message exchange between atoms in reciprocal space Kosmala et al. (2023).
III Examples
III.1 Molecular dimers
@@
We benchmark the LES method on the binding curves between dimers of charged (C), polar (P), and apolar (A) relaxed molecules at various separations in a periodic cubic box with a 30 Å edge length. The dataset Huguenin-Dumittan et al. (2023) includes energy and force information calculated using the HSE06 hybrid density functional theory (DFT) with a many-body dispersion correction. We selected one example from each of the six dimer classes (CC, CP, PP, CA, CP, AA), derived from the combination of the three monomer categories. Figure 1 shows a snapshot of each example.
@@
The goal of this benchmark is to evaluate whether the MLIP models can extrapolate dimer interactions at larger separations based on training data from smaller distances. For each molecular pair, the training set consists of 10 configurations with dimer separation distances between approximately 5 Å and 12 Å, and the testing set includes 3 configurations with separations between about 12 Å and 15 Å.
@@
Figure 1: Comparison of the short-range (SR) and long-range (LR) machine learning interatomic potential performance for six dimer classes: charged-charged (CC), charged-polar (CP), polar-polar (PP), charged-apolar (CA), polar-apolar (PA), apolar-apolar (AA). For each class, the upper panel shows a snapshot of the system with the charge states indicated, the middle panel shows the parity plot for the force components, and the lower panel shows the binding energy curve, which is the potential energy difference between the dimer and two isolated monomers. The root mean square errors (RMSE) for the energy and force components of the test sets are shown in the insets.
@@
For each molecular pair, we trained a short-range (SR) model using CACE with a cutoff $r_{\mathrm{cut}}=5$, 6 Bessel radial functions, $c=8$, $l_{\mathrm{max}}=2$, $\nu_{\mathrm{max}}=2$, $N_{\mathrm{embedding}}=3$, and one message passing layer ($T=1$). It is worth noting that this “SR” model setting already achieves a perceptive field of 10 Å via the message passing layer, which is quite typical for current MPNNs. In comparison, more traditional MLIPs based on local atomic descriptors typically use a cutoff of around 5 or 6 Å, making them even more short-ranged.
@@
Fig. 1 demonstrates that the LR MLIPs outperform the SR MLIPs across all cases. The improvement is particularly pronounced for the CC, CP, and PP cases, where the root mean square errors (RMSEs) of energy and force components can be reduced by an order of magnitude. Additionally, the SR models fail to adequately capture the training data, as indicated by the flattening of the binding curves and the large discrepancies between some predicted and true forces (gray symbols in Fig. 1). The primary limitation of the SR models is that molecules separated by distances beyond the MLIP cutoff exist on independent atomic graphs, rendering the message passing layers ineffective for communication between the dimers. The improvement achieved by the LR models for the CA, PA, and AA cases is more modest, though the energy scales in these systems are much smaller, and the absolute RMSE values are relatively low. The residual RMSEs can also be attributed to the limited (only 10) training data points used in each case. Direct comparison of the accuracy of the present LR models with the LODE method Huguenin-Dumittan et al. (2023) is challenging, as different training protocols were employed in Ref. Huguenin-Dumittan et al. (2023), and LODE depends significantly on the choice of the potential exponent $p$ for each dimer class. However, based solely on RMSE values, the LR models presented here appear to be more accurate. For example, in the CC class, the energy RMSE is 15.5 meV, compared to approximately 0.1 eV for LODE with the optimal choice of $p$.
III.2 Bulk water
@@
As an example application to dipolar fluids, we applied the LES method to a dataset of 1,593 liquid water configurations, each containing 64 molecules Cheng et al. (2019). The dataset was computed using revPBE0-D3 DFT. For the CACE representation, we employed a cutoff of $r_{\mathrm{cut}}=5.5$, 6 Bessel radial functions with $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, and either no message passing ($T=0$) or one message passing layer ($T=1$). For the long-range component, we used a 4-dimensional $q$, a maximum cutoff of $k_{c}=\pi$, and $\sigma=1$ in the Ewald summation.
@@
The learning curves in Fig. 2 demonstrate that message passing significantly improves the accuracy of the MLIPs, consistent with previous studies Batzner et al. (2022); Batatia et al. (2022). The LR component further reduces the error for both models without a message passing layer ($T=0$) and with one message passing layer ($T=1$). The improvement is particularly notable in the $T=0$ scenario. For the $T=1$ models, the LR component results in a smaller reduction in errors, likely because the $T=1$ SR models already capture atomic interactions up to 11 Å. Nevertheless, the $T=1$ LR model demonstrates greater efficiency in learning with fewer data. We also compared the inference speeds of these models during molecular dynamics (MD) simulations, as detailed in the Methods section.
@@
To investigate how message passing and long-range interactions affect predicted structural properties, we performed MD simulations of bulk water with a density of 1 g/mL at 300 K using each model. The upper panel of Fig. 3 shows the oxygen-oxygen (O-O) radial distribution function (RDF) computed with different MLIPs. All the computed O-O RDFs are indistinguishable and in excellent agreement with experimental results from X-ray diffraction measurements Skinner et al. (2014). This suggests that local representations are sufficient for accurately predicting the RDFs of bulk liquid water, consistent with previous findings Yue et al. (2021).
@@
As shown in Fig. 4, while the dipole density correlation functions predicted by different MLIPs are in excellent agreement for most values of $k$, discrepancies emerge at long wavelengths. Specifically, the results from short-ranged models sharply increase as $k\rightarrow 0$. This divergence has also been observed in previous studies comparing SR and LR models Cox (2020); Gao and Remsing (2022). The divergence of the $T=0$ short-ranged MLIP appears at a $k$ value corresponding to a real length scale of about 11 Å, while the $T=1$ SR model diverges at about 16 Å. Interestingly, these length scales, where divergence occurs, exceed the effective cutoff radius, suggesting that SR MLIPs may partially describe long-range effects beyond their cutoffs in a mean-field manner. Comparing the two SR MLIPs, the message passing layer delays the onset of divergence at small $k$ values but does not eliminate it. In this scenario, only true LR models can adequately describe the dielectric response.

### v2

@@ Abstract
Machine learning interatomic potentials (MLIPs) often neglect long-range interactions, such as electrostatic and dispersion forces. In this work, we introduce a straightforward and efficient method to account for long-range interactions by learning a hidden variable from local atomic descriptors and applying an Ewald summation to this variable. We demonstrate that in systems including charged and polar molecular dimers, bulk water, and water-vapor interface, standard short-ranged MLIPs can lead to unphysical predictions even when employing message passing. The long-range models effectively eliminate these artifacts, with only about twice the computational cost of short-range MLIPs.
@@
Machine learning interatomic potentials (MLIPs) can learn from reference quantum mechanical calculations and then predict the energy and forces of atomic configurations quickly, allowing for a more accurate and comprehensive exploration of material and molecular properties at scale Keith et al. (2021); Unke et al. (2021a). Most state-of-the-art MLIP methods use a short-range approximation: the effective potential energy surface experienced by one atom is determined by its atomic neighborhood. This approximation implies that the total energy is the sum of atomic contributions, which also makes the MLIPs scale linearly with system size.
@@
The short-range MLIPs, however, neglect all kinds of long-range interactions such as Coulomb and dispersion. Although short-range potentials may be sufficient to describe most properties of homogeneous bulk systems Yue et al. (2021), they may fail for liquid-vapor interfaces Niblett et al. (2021), dielectric response Rodgers and Weeks (2008); Cox (2020), dilute ionic solutions with Debye-Hückel screening, and interactions between gas phase molecules Huguenin-Dumittan et al. (2023).
@@
There has been a continuous effort to incorporate long-range interactions into MLIPs. One can include empirical electrostatics and dispersion baseline corrections Niblett et al. (2021); Lee et al. (2022); Unke et al. (2021b), but for many systems, such baseline is not readily available. Another option is to predict effective partial charges to each atom, which are then used to calculate long-range electrostatics Unke and Meuwly (2019); Ko et al. (2021); Gao and Remsing (2022); Sifain et al. (2018); Gong et al. (2024); Shaidu et al. (2024). For example, the fourth-generation high-dimensional neural network potential (4G-HDNNPs) Ko et al. (2021) predicts the electronegativities of each nucleus, and then use a charge equilibration scheme Rappe and Goddard III (1991) to assign the charges. 4G-HDNNPs are trained directly to reproduce atomic partial charges from reference quantum mechanical calculations, although partial charges are not physical observable and their values depend on the specific partitioning scheme used Sifain et al. (2018). In a similar vein, the deep potential long-range (DPLR) Zhang et al. (2022) learns maximally localized Wannier function centers (MLWFCs) for insulating systems, and the self-consistent field neural network (SCFNN) Gao and Remsing (2022) predicts the electronic response via the position of the MLWFCs. Message passing neural networks (MPNNs) Schütt et al. (2017); Batzner et al. (2022); Deng et al. (2023); Haghighatlari et al. (2022) employ a number of graph convolution layers to communicate information between atoms, thus capturing long-range interaction up to the local cutoff radius times the number of layers. However, if parts of the system are disconnected on the graph, e.g. two molecules with a distance beyond the cutoff, the message passing scheme does not help. Another class of methods is to learn the long-range descriptors and interactions in the reciprocal space with learnable frequency filters Yu et al. (2022); Kosmala et al. (2023). Finally, a very interesting approach is the long-distance equivariant (LODE) method Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023), which uses local descriptors to encode the Coulomb and other asymptotic decaying potentials ($1/r^{p}$) around the atoms, and a related, density-based long-range descriptors Faller et al. (2024).
@@ I.1 Theory
where $E_{\theta}$ is a multilayer perceptron with parameters $\theta$ that maps the invariant features ($B$) of an atom to its short-range atomic energy. $B$ can be any invariant features used in different MLIP methods, including those based on local atomic environment descriptors such as ACE Drautz (2019), atom-centered symmetry functions Behler and Parrinello (2007), smooth overlap of atomic positions (SOAP) Bartók et al. (2013), or any latent invariant features in MPNNs.
@@
where $\mathbf{k}=(2\pi n_{x}/L_{x},2\pi n_{y}/L_{y},2\pi n_{z}/L_{z})$ is a reciprocal vector of the orthorhombic cell, and $\mathbf{r}_{i}$ is the Cartesian coordinates of atom $i$. The long-range energy is then obtained using an Ewald summation form that best captures the electrostatic potential ($1/r$Williams (1971) :
@@
where $\sigma$ is a smearing factor which we typically set to 1 Å with justifications in the Methods , $k=|\mathbf{k}|$ is the magnitude, and $k_{c}$ is the maximum cutoff. $q$ can be multi-dimensional, in which case the total long-range energy is aggregated over contributions from different dimensions of $q$ after the Ewald summation.
@@
The LES method can be interpreted in two ways. First, the hidden variable $q$ is analogous to the environmental-dependent partial charges on each atom. This implies that the method is at least as expressive as those explicitly based on learning partial atomic charges, as it would yield the same results if $q$ replicated the partial charges. Additionally, $q$ can be multidimensional, potentially enhancing expressiveness further. Unlike partial charges, $q$ is not constrained by requirements such as charge neutrality or correct charge magnitudes. As the Ewald summation in Eqn. (4) omits the $k=0$ term, a non-zero net $q$ does not cause energy divergence issues. Since the Ewald summation of $q$ doesn’t need to correspond to a physical electrostatic potential, Eqn. (4) omits the self-interaction term present in the regular Ewald summation for long-range charge interactions. The second interpretation of LES is as a mechanism that allows atoms far apart in the simulation box to communicate their local information. In this sense, LES is related to the recent Ewald-based long-range message passing method, which facilitates message exchange between atoms in reciprocal space Kosmala et al. (2023).
I.2 Example on molecular dimers
@@
We benchmark the LES method on the binding curves between dimers of charged (C) and polar (P) molecules at various separations in a periodic cubic box with a 30 Å edge length. The dataset Huguenin-Dumittan et al. (2023), originally from the BioFragment Database Burns et al. (2017), includes energy and force information calculated using the HSE06 hybrid density functional theory (DFT) with a many-body dispersion correction. We selected one example from each of the three dimer classes (CC, CP, PP) , derived from the combination of the three monomer categories. Figure 1 shows a snapshot of each example.
@@
The goal of this benchmark is to evaluate whether the MLIP models can extrapolate dimer interactions at larger separations based on training data from smaller distances. For each molecular pair, the training set consists of 10 configurations with dimer separation distances between approximately 5 Å and 12 Å, and the test set includes 3 configurations with separations between approximately 12 Å and 15 Å.
@@
Figure 1: Comparison of the short-range (SR) and long-range (LR) machine learning interatomic potential performance for three dimer classes: charged-charged (CC), charged-polar (CP), and polar-polar (PP). For each class, the upper panel shows a snapshot of the system with the charge states indicated, the middle panel shows the parity plot for the force components, and the lower panel shows the binding energy curve, which is the potential energy difference between the dimer, and two isolated and relaxed monomers. The root mean square errors (RMSE) for the energy and force components of the test sets are shown in the insets.
@@
For each molecular pair, we trained a short-range (SR) model using CACE with a cutoff $r_{\mathrm{cut}}=5$, 6 Bessel radial functions, $c=8$, $l_{\mathrm{max}}=2$, $\nu_{\mathrm{max}}=2$, $N_{\mathrm{embedding}}=3$, and one message passing layer ($T=1$). It should be noted that this setting of the “SR” model already achieves a perceptive field of 10 Å through the message passing layer, which is quite typical for current MPNNs. In comparison, more traditional MLIPs based on local atomic descriptors typically use a cutoff of around 5 or 6 Å, making them even more short-ranged.
@@
Fig. 1 shows that the LR MLIPs outperform the SR MLIPs in all cases. Furthermore, SR models fail to adequately capture the training data, as indicated by flattening of the binding curves and the large discrepancies between some predicted and true forces (gray symbols in Fig. 1). The primary limitation of the SR models is that molecules separated by distances beyond the MLIP cutoff exist on independent atomic graphs, rendering the message-passing layers ineffective for communication between the dimers. Direct comparison of the accuracy of the present LR models with the LODE method Huguenin-Dumittan et al. (2023) is challenging, as different training protocols were used in Ref. Huguenin-Dumittan et al. (2023), and LODE depends significantly on the choice of the potential exponent $p$ for each dimer class. However, based solely on RMSE values, the LR models presented here appear to be more accurate. For example, in the CC class, the energy RMSE is 15.5 meV, compared to approximately 0.1 eV for LODE with the optimal choice of $p$.
I.3 Example on molten NaCl
Molten bulk sodium chloride (NaCl) presents non-negligible long-range electrostatic interactions. The dataset from Ref. Faller et al. (2024) contains 1014 structures (80% train and 20% validation) of 64 Na and 64 Cl atoms. Table 1 compares the root-mean-square percentage error (RMSPE) using different methods. $r_{\mathrm{cut}}=~{}$6 Å were used for all models except for the LODE flexible (7 Å). The CACE-LR models use 6 Bessel radial functions, $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, zero ($T=0$) or one message passing layer ($T=1$), a 4-dimensional $q$, $\sigma=1$ and a $\mathbf{k}$-point cutoff of $k_{c}=2\pi/3$ in Ewald summation. The details for the other models are in Ref. Faller et al. (2024).
Table 1: The root-mean-square percentage error (RMSPE) for the validation set of molten NaCl, using the short-ranged SOAP descriptor Bartók et al. (2010, 2013); Jinnouchi et al. (2019), two setting of LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) descriptors (minimal and flexible), MACE Batatia et al. (2022) with zero ($T=0$) or one message passing layers ($T=1$) as well as using invariant and equivariant messages, a density-based long-range model (Density-LR) Faller et al. (2024), and CACE-LR with zero ($T=0$) or one message passing layers ($T=1$). The fits other than the CACE models are from Ref. Faller et al. (2024).
RMSPE E (%) RMSPE F (%)
SOAP Bartók et al. (2010, 2013); Jinnouchi et al. (2019) 7.7 12.6
MACE Batatia et al. (2022) T=0 8.5 8.5
MACE Batatia et al. (2022) T=1 invariant 3.0 3.1
MACE Batatia et al. (2022) T=1 equivariant 2.1 2.2
LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) minimal 5.6 3.7
LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) flexible 7.0 11.0
Density-LR Faller et al. (2024) 2.2 3.3
CACE-LR T=0 1.9 2.3
CACE-LR T=1 1.4 1.8
Table 1 shows that the short-ranged models, including SOAP Bartók et al. (2010, 2013); Jinnouchi et al. (2019) and MACE Batatia et al. (2022) $T=0$, have the worst performances. All the long-ranged models, including LODE Grisafi and Ceriotti (2019); Huguenin-Dumittan et al. (2023) with different settings, a density-based long-range model Faller et al. (2024), and CACE-LR outperform the purely SR methods. MACE Batatia et al. (2022) with message passing ($T=1$), especially the one using equivariant features, also achieves accurate results, as message passing increases the effective perceptive field. Overall, CACE-LR models, with and without message passing ($T=0$ and $T=1$), obtain the best accuracy.
I.4 Example on bulk water
@@
As an example application to dipolar fluids, we applied the LES method to a dataset of 1,593 liquid water configurations, each containing 64 molecules Cheng et al. (2019). The dataset was calculated using revPBE0-D3 DFT. For the CACE representation, we used a cutoff of $r_{\mathrm{cut}}=5.5$, 6 Bessel radial functions with $c=12$, $l_{\mathrm{max}}=3$, $\nu_{\mathrm{max}}=3$, $N_{\mathrm{embedding}}=3$, and no message passing ($T=0$) or no message passing layer ($T=1$). For the long-range component, we used a 4-dimensional $q$, a maximum cutoff of $k_{c}=\pi$, and $\sigma=1$ in the Ewald summation.
@@
The learning curves in Fig. 2 demonstrate that message passing significantly improves the accuracy of MLIPs, consistent with previous studies Batzner et al. (2022); Batatia et al. (2022). The LR component further reduces the error for both models without a message passing layer ($T=0$) and with a message passing layer ($T=1$). The improvement is particularly notable in the $T=0$ scenario. For the $T=1$ models, the LR component results in a smaller reduction in errors, probably because the $T=1$ SR models already capture atomic interactions up to 11 Å. Nevertheless, the $T=1$ LR model demonstrates greater efficiency in learning with fewer data. We also compared the inference speeds of these models during molecular dynamics (MD) simulations, as detailed in the Methods section.
@@
To investigate how message passing and long-range interactions affect predicted structural properties, we performed MD simulations of bulk water with a density of 1 g/mL at 300 K using each model. The upper panel of Fig. 3 shows the oxygen-oxygen (O-O) radial distribution function (RDF) computed with different MLIPs. All computed O-O RDFs are indistinguishable and in excellent agreement with the experimental results of the X-ray diffraction measurements Skinner et al. (2014). This suggests that local representations are sufficient to accurately predict the RDFs of bulk liquid water, consistent with previous findings Yue et al. (2021).
@@
As shown in Fig. 4, while the dipole density correlation functions predicted by different MLIPs are in excellent agreement for most values of $k$, discrepancies emerge at long wavelengths. Specifically, the results from short-ranged models sharply increase as $k\rightarrow 0$. This divergence has also been observed in previous studies comparing the SR and LR models Cox (2020); Gao and Remsing (2022). The divergence of the $T=0$ short-ranged MLIP appears at a $k$ value corresponding to a real length scale of approximately 11 Å, while the $T=1$ SR model diverges at about 16 Å. Interestingly, these length scales, where divergence occurs, exceed the effective cutoff radius, suggesting that SR MLIPs may partially describe long-range effects beyond their cutoffs in a mean-field manner. When comparing the two SR MLIPs, the message passing layer delays the onset of divergence at small $k$ values but does not eliminate it. In this scenario, only true LR models can adequately describe the dielectric response.